# Contrastive Learning on Multimodal Analysis of Electronic Health Records

### v1

@@ 3 Theoretical Analysis
The proof of Theorem 3.2 is given in Supplementary S4.3. By Assumption 3.2, it implies that $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$ is an asymptotically unbiased estimate of $\mathbb{PMI}^{(1,2)}$. Combining Theorems 3.1 and 3.2, we get the rate of CLAIME. We also provide a rate of learning dual representations jointly by CLAIME. Namely, we consider concatenating the left singular vectors of two matrices $\mathbf{\widehat{V}}_{1}$ and $\mathbf{\widehat{V}}_{2}$. Since $\mathbf{P}_{p}(\mathbf{\widehat{V}}_{1}^{\top}),\mathbf{P}_{p}(\mathbf{\widehat{V}}_{2}^{\top})$ are only identifiable up to rotation, the subspace spanned by the column vectors of $\mathbf{P}_{p}([\mathbf{P}_{p}^{\top}(\mathbf{\widehat{V}}_{1}^{\top}),\mathbf{P}_{p}^{\top}(\mathbf{\widehat{V}}_{2})]$ depends on the choice of rotations. Hence we introduce a matrix $\mathbf{H}$, and define the estimated joint representation by $\mathbf{\widehat{U}}_{\mathbf{H}}:=\mathbf{P}_{p}([\mathbf{H}\mathbf{P}_{p}^{\top}(\mathbf{\widehat{V}}_{1}^{\top}),\mathbf{P}_{p}^{\top}(\mathbf{\widehat{V}}_{2}^{\top})])$, where $\mathbf{P}_{p}(\mathbf{\widehat{V}}_{1}^{\top})$ and $\mathbf{P}_{p}(\mathbf{\widehat{V}}_{2}^{\top})$ are arbitrary top-$p$ left singular vectors of $\mathbf{\widehat{V}}_{1}$ and $\mathbf{\widehat{V}}_{2}$, respectively. We then measure the best performance of $\mathbf{\widehat{U}}_{\mathbf{H}}$ with respect to $\mathbf{H}$ by $\min_{\mathbf{H}\in\mathcal{O}_{p,p}}\|\sin\Theta(\mathbf{\widehat{U}}_{\mathbf{H}},\mathbf{U}^{\star})\|_{\mathrm{F}}$, where $\mathbf{U}^{\star}\in\mathcal{O}_{d,p}$ is the concatenated embedding vectors defined as $\mathbf{U}^{\star}:=\mathbf{P}_{p}([\mathbf{U}_{1}^{\star\top},\mathbf{U}_{2}^{\star\top}])$.
@@
Suppose that $n\gg d^{2}\log^{2}d$. Under Assumptions 3.1-3.3 and 3.5,
$\norm{\mathbf{\widehat{V}}_{1}\mathbf{\widehat{V}}_{2}^{\top}-\frac{1}{\lambda p}\mathbf{V}_{1}^{\star}\mathbf{V}_{2}^{\star\top}}\lesssim\frac{1}{\lambda}\quantity(\frac{\underline{d}^{3/2}d}{\sqrt{n}}\log d+\frac{p^{3/2}}{\underline{d}}\log^{6}d),$
$\|\sin\Theta(\mathbf{P}_{p}(\mathbf{\widehat{V}}_{1}^{\top}),\mathbf{U}_{1}^{\star})\|_{\mathrm{F}}\vee\|\sin\Theta(\mathbf{P}_{p}(\mathbf{\widehat{V}}_{2}^{\top}),\mathbf{U}_{2}^{\star})\|_{\mathrm{F}}\lesssim\frac{p^{3/2}\underline{d}^{3/2}d}{\sqrt{n}}\log d+\frac{p^{3}}{\underline{d}}\log^{6}d$
and
$\min_{\mathbf{H}\in\mathcal{O}_{p,p}}\|\sin\Theta(\mathbf{\widehat{U}}_{\mathbf{H}},\mathbf{U}^{\star})\|_{\mathrm{F}}$ $\lesssim\frac{p^{3/2}\underline{d}^{3/2}d}{\sqrt{n}}\log d+\frac{p^{3}}{\underline{d}}\log^{6}d$
with probability $1-\exp(-\Omega(\log^{2}d))$.
The proof of Theorem 3.3 is presented in Supplementary S4.3. On the contrary, we can show that the empirical concatenated PMI matrix $\widehat{\mathbb{PMI}}$ and $\widehat{\mathbb{PMI}}_{\textnormal{CL}}$ derived from the CL loss (2.6) are biased estimators for $\mathbb{PMI}$.
@@
Suppose that $n\gg d^{2}\log^{2}d$. Under Assumptions 3.1-3.3 and 3.5, we have
$\norm{\widehat{\mathbb{PMI}}-(\mathbb{PMI}+\mathbf{B}_{\textnormal{Con}})}_{\max}\lesssim\frac{d^{2}}{\sqrt{n}}\log d,\quad\norm{\widehat{\mathbb{PMI}}-(\mathbb{PMI}+\mathbf{B}_{\textnormal{Con}})}\lesssim\frac{d^{5/2}}{\sqrt{n}}\log d,$
and
$\norm{\widehat{\mathbb{PMI}}_{\textnormal{CL}}-(\mathbb{PMI}+\mathbf{B}_{\textnormal{CL}})}_{\max}$ $\lesssim\frac{d^{2}}{\sqrt{n}}\log d+\frac{p^{3/2}}{\underline{d}^{3/2}}\log^{6}d,$
$\norm{\widehat{\mathbb{PMI}}_{\textnormal{CL}}-(\mathbb{PMI}+\mathbf{B}_{\textnormal{CL}})}$ $\lesssim\frac{d^{5/2}}{\sqrt{n}}\log d+\quantity(\frac{d}{\underline{d}})^{1/2}\frac{p^{3/2}}{\underline{d}}\log^{6}d.$
Here the matrices $\mathbf{B}_{\textnormal{Con}}$ and $\mathbf{B}_{\textnormal{CL}}$ reflect the bias terms defined in Supplementary S4.2 and S4.4.
Remark 3.1.
The bias matrices $\mathbf{B}_{\textnormal{Con}}$ and $\mathbf{B}_{\textnormal{CL}}$ depend only on $T_{i}^{(M)}$’s, which can be effectively removed since $T_{i}^{(M)}$’s are directly observed.
The proof of Theorem 3.4 is presented in Supplementary S4.4. Based on this theorem, we are ready to show the rate of $\widehat{\mathbf{V}}_{\textnormal{CL}}$ and $\mathbf{\widehat{V}}_{\textnormal{Con}}$. Before that, we need the following assumption.
@@
Assume that
$\frac{s_{p}({\bm{\Sigma}})-s_{p+1}({\bm{\Sigma}})}{s_{p}({\bm{\Sigma}})}\gtrsim 1.$
Assumption 3.4 is known as the eigengap condition for the top-$p$ singular vectors of ${\bm{\Sigma}}$. A similar assumption has been used in the analysis of contrastive learning (Ji et al.,, 2021). This enables us to identify the top-$p$ singular vectors of $(1/p)\mathbf{V}^{\star}\mathbf{V}^{\star\top}+{\bm{\Sigma}}$, resulting in the following theorem.
Theorem 3.5.
Suppose that $n\gg d^{2}\log^{2}d$ and $\|\mathbf{B}_{\textnormal{Con}}\|\vee\|\mathbf{B}_{\textnormal{CL}}\|\ll 1/p$. Under Assumptions 3.1-3.5,
$\norm{\sin\Theta\quantity(\mathbf{P}_{p}(\mathbf{\widehat{V}}_{\textnormal{Con}}^{\top}),\mathbf{P}_{p}\quantity(\frac{1}{p}\mathbf{V}^{\star}\mathbf{V}^{\star\top}+{\bm{\Sigma}}+\mathbf{B}_{\textnormal{Con}}))}_{\mathrm{F}}$ $\lesssim\frac{p^{3/2}d^{5/2}}{\sqrt{n}}\log d+\frac{p^{3}}{\underline{d}}\log^{6}d,$
$\norm{\sin\Theta\quantity(\mathbf{P}_{p}(\widehat{\mathbf{V}}_{\textnormal{CL}}^{\top}),\mathbf{P}_{p}\quantity(\frac{1}{p}\mathbf{V}^{\star}\mathbf{V}^{\star\top}+{\bm{\Sigma}}+\mathbf{B}_{\textnormal{CL}}))}_{\mathrm{F}}$ $\lesssim\frac{p^{3}d^{5/2}}{\sqrt{n}}\log d+\quantity(\frac{d}{\underline{d}})^{1/2}\frac{p^{9/2}}{\underline{d}}\log^{6}d$
holds with probability $1-\exp(-\Omega(\log^{2}d))$.
The proof Theorem 3.5 is given in Supplementary S4.4. It shows that both Concate and CL will yield biased estimators for the embeddings $\mathbf{V}^{\star}$ due to the inherent bias terms and covariance matrix of the noise. While the bias terms $\mathbf{B}_{\textnormal{Con}}$ and $\mathbf{B}_{\textnormal{CL}}$ can be removed, the influence of ${\bm{\Sigma}}$ remains irreducible due to its unobserved nature. Comparing Theorems 3.3 and 3.5, we find that CLAIME achieves a sharper rate than Concate and CL even when $\mathbf{B}_{\textnormal{CL}}$, $\mathbf{B}_{\textnormal{Con}}$ and ${\bm{\Sigma}}$ go to zero. As a result, CLAIME is a better estimator than Concate and CL.
To quantify the effect of $\mathbf{B}_{\textnormal{Con}}$ and $\mathbf{B}_{\textnormal{CL}}$, we define the following empirical moments of $T_{i}^{(M)}$: $S_{q}^{(M)}:=((1/n)\sum_{i\in[n]}T_{i}^{(M)q})^{1/q}$. We also define the empirical cross-moments as $S_{q}^{(1,2)}:=((1/n)\sum_{i\in[n]}(T_{i}^{(1)})^{q/2}(T_{i}^{(2)})^{q/2})^{1/q}$ for $q=2,4$. Note that when the length of patient data is the same, i.e., $T_{i}^{(M)}=T^{(M)}$ for all $i\in[n]$ for $M=1,2$, $S_{q}^{(M)}=T^{(M)}$ for all $q\geq 1$.
Assumption 3.5.
For $M=1,2$, assume that $1\gtrsim S_{1}^{(1)}S_{1}^{(2)}/S_{2}^{(1,2)2}\gg 1/n$, $S_{1}^{(M)2}/S_{2}^{(M)2}\gg 1/n$ and $S_{2}^{(M)2}\gg S_{1}^{(M)}$. Also assume that $S_{1}^{(M)}\gtrsim S_{2}^{(M)}\gtrsim S_{4}^{(M)}$ and $S_{2}^{(1,2)}\gtrsim S_{4}^{(1,2)}$.
Assumption 3.5 requires that the empirical variance of $T_{i}^{(M)}$ dominates the mean of $T_{i}^{(M)}$, and their moments are comparable to each other. For the same length setting $T_{i}^{(M)}\equiv T^{(M)}$ for all $i\in[n]$, the assumption is satisfied if $T^{(M)}\to\infty$ and $n\to\infty$. For the bias term $\mathbf{B}_{\textnormal{Con}}$ and $\mathbf{B}_{\textnormal{CL}}$, we have the following lemma. The proof of Lemma 3.1 is given in Supplementary S4.7.
Lemma 3.1.
Suppose that Assumption 3.5 holds. If $T_{i}^{(1)}=T_{i}^{(2)}$, then,
$\|\mathbf{B}_{\textnormal{Con}}\|\vee\|\mathbf{B}_{\textnormal{CL}}\|\lesssim d\frac{S_{1}^{(1)}}{S_{2}^{(1)2}}.$
Furthermore, if $T_{i}^{(1)}=T_{i}^{(2)}\equiv T$ for all $i\in[n]$ for some $T>0$, then
$\|\mathbf{B}_{\textnormal{Con}}\|\vee\|\mathbf{B}_{\textnormal{CL}}\|\lesssim\frac{d}{T}.$
We then characterize the effect of ${\bm{\Sigma}}$. We will show that when ${\bm{\Sigma}}$ satisfies suitable conditions, the two existing methods Concate and CL will not be an ideal estimator of $\mathbf{V}^{\star}$.
@@
Suppose that Assumptions 3.1-3.6 hold. If $\|\mathbf{B}_{\textnormal{Con}}\|\vee\|\mathbf{B}_{\textnormal{CL}}\|\ll 1$, then,
$\|\sin\Theta(\mathbf{\widehat{U}}_{\textnormal{Con}},\mathbf{U}^{\star})\|_{\mathrm{F}}$ $\gtrsim\sqrt{p}\ \ \text{ and }\ \ \|\sin\Theta(\mathbf{\widehat{U}}_{\textnormal{CL}},\mathbf{U}^{\star})\|_{\mathrm{F}}\gtrsim\sqrt{p}$
with probability $1-\exp(-\Omega(\log^{2}d))$.
@@
Choose $\mathbf{P}^{(1)}\in\mathcal{O}_{d_{1},p}$ satisfying $[\mathbf{U}_{1}^{\star};\mathbf{P}^{(1)}]\in\mathcal{O}_{d_{1},2p}$ and $\|[\mathbf{U}_{1}^{\star};\mathbf{P}^{(1)}]\|_{2,\infty}^{2}\lesssim p/d_{1}$. Similarly choose $\mathbf{P}^{(2)}\in\mathcal{O}_{d_{2},p}$ such that $[\mathbf{U}_{2}^{\star};\mathbf{P}^{(2)}]\in\mathcal{O}_{d_{2},2p}$ and $\|[\mathbf{U}_{2}^{\star};\mathbf{P}^{(2)}]\|_{2,\infty}^{2}\lesssim p/d_{2}$. Let ${\bm{\Sigma}}_{1}:=\mathbf{P}^{(1)}\mathbf{P}^{(1)\top}$, ${\bm{\Sigma}}_{2}:=(1/2)\mathbf{P}^{(2)}\mathbf{P}^{(2)\top}$, $\mathbf{V}^{\star}_{1}=\mathbf{U}^{\star}_{1}$ and $\mathbf{V}^{\star}_{2}=\mathbf{U}^{\star}_{2}$. For these noise covariance matrices, Assumptions 3.1, 3.3 and 3.4 hold. In addition, Assumption 3.6 holds since $\mathbf{P}^{(1)\top}\mathbf{V}_{1}^{\star}=\mathbf{P}^{(2)\top}\mathbf{V}_{2}^{\star}=\mathbf{O}_{p\times p}$. Under additional assumptions on $(T_{i}^{(M)})_{i\in[n]}$, $d_{1}$ and $d_{2}$ (Assumptions 3.2 and 3.5), we obtain
@@
$\min_{\mathbf{H}\in\mathcal{O}_{p,p}}\|\sin\Theta(\mathbf{\widehat{U}}_{\mathbf{H}},\mathbf{U}^{\star})\|_{\mathrm{F}}$ $\lesssim\frac{p^{3/2}\underline{d}^{3/2}d}{\sqrt{n}}\log d+\frac{p^{3}}{\underline{d}}\log^{6}d.$
Corollary 3.1 directly follows from Theorems 3.3 and 3.6. Note that the factor $1/2$ in the definition of ${\bm{\Sigma}}_{2}$ guarantees the eigengap condition at the $p$-th largest singular value of the joint noise covariance matrix ${\bm{\Sigma}}$. The noise covariance in Corollary 3.1 implies that the subspace spanned by the noise is orthogonal to the subspace spanned by the signal, so that the noisy fluctuations in feature frequency distribution occurs simultaneously to many words, in a separable way from the signal. Under this condition, Concate and CL learn the eigenspace corresponding to noise covariance matrix, since those methods are vulnerable to large noise strength. On the other hand, CLAIME can effectively escape from the noise, and learns core representations when $n$ and $\bar{d}$ are sufficiently large.

### v2

@@ 3.1 Theoretical Properties of CLAIME
The proof of Theorem 3.2 is given in Supplementary S4.3. Under Assumption 3.2, $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$ is an asymptotically unbiased estimate of $\mathbb{PMI}^{(1,2)}$. Combining Theorems 3.1 and 3.2 yields the convergence rate of CLAIME. We also analyze the rate of learning joint representations via CLAIME. Specifically, we concatenate the left singular vectors of $\mathbf{\widehat{V}}_{1}$ and $\mathbf{\widehat{V}}_{2}$. Since $\mathbf{P}_{p}(\mathbf{\widehat{V}}_{1}^{\scriptscriptstyle\sf T}),\mathbf{P}_{p}(\mathbf{\widehat{V}}_{2}^{\scriptscriptstyle\sf T})$ are only identifiable up to rotation, the subspace spanned by $\mathbf{P}_{p}([\mathbf{P}_{p}^{\scriptscriptstyle\sf T}(\mathbf{\widehat{V}}_{1}^{\scriptscriptstyle\sf T}),\mathbf{P}_{p}^{\scriptscriptstyle\sf T}(\mathbf{\widehat{V}}_{2})]$ depends on the choice of rotations. To address this, we introduce a matrix $\mathbf{H}$ and define the estimated joint representation by $\mathbf{\widehat{U}}_{\mathbf{H}}:=\mathbf{P}_{p}([\mathbf{H}\mathbf{P}_{p}^{\scriptscriptstyle\sf T}(\mathbf{\widehat{V}}_{1}^{\scriptscriptstyle\sf T}),\mathbf{P}_{p}^{\scriptscriptstyle\sf T}(\mathbf{\widehat{V}}_{2}^{\scriptscriptstyle\sf T})])$. We then evaluate performance via $\min_{\mathbf{H}\in\mathcal{O}_{p,p}}\|\sin\Theta(\mathbf{\widehat{U}}_{\mathbf{H}},\mathbf{U}^{\star})\|_{\mathrm{F}}$, where $\mathbf{U}^{\star}\in\mathcal{O}_{d,p}$ is the concatenated embedding vectors defined as $\mathbf{U}^{\star}:=(1/\sqrt{2})[\mathbf{U}_{1}^{\star{\scriptscriptstyle\sf T}},\mathbf{U}_{2}^{\star{\scriptscriptstyle\sf T}}]$. The proof of Theorem 3.3 is presented in Supplementary S4.3.
Theorem $3.$
Under Assumptions 3.1-3.4, with probability $1-\exp(-\Omega(\log^{2}(n+d)))$,
$\|\sin\Theta(\mathbf{P}_{p}(\mathbf{\widehat{V}}_{1}^{\scriptscriptstyle\sf T}),\mathbf{U}_{1}^{\star})\|_{\mathrm{F}}\vee\|\sin\Theta(\mathbf{P}_{p}(\mathbf{\widehat{V}}_{2}^{\scriptscriptstyle\sf T}),\mathbf{U}_{2}^{\star})\|_{\mathrm{F}}\lesssim\frac{p^{3}(d_{1}\vee d_{2})\log(n+d)}{(d_{1}\wedge d_{2})^{2}}$
$\quad\quad+\frac{pd_{1}^{3/2}d_{2}^{3/2}\log(n+d)}{\sqrt{n}(S_{1}^{(1)1/2}\wedge S_{1}^{(2)1/2})}+\frac{pd_{1}^{1/2}d_{2}^{1/2}\log(n+d)}{\sqrt{n}}\quad\text{ and }$ (3.3)
$\min_{\mathbf{H}\in\mathcal{O}_{p,p}}\|\sin\Theta(\mathbf{\widehat{U}}_{\mathbf{H}},\mathbf{U}^{\star})\|_{\mathrm{F}}$ $\lesssim\left\{\frac{p^{3}(d_{1}\vee d_{2})}{(d_{1}\wedge d_{2})^{2}}+\frac{pd_{1}^{3/2}d_{2}^{3/2}}{\sqrt{n}(S_{1}^{(1)1/2}\wedge S_{1}^{(2)1/2})}+\frac{pd_{1}^{1/2}d_{2}^{1/2}}{\sqrt{n}}\right\}\log(n+d).$
Remark 3.1.
The bound in (3.3) consists of a bias term from the nonlinearity of the data-generating process (first term), and two variance terms (last two terms). The first variance term arises from estimating per-record co-occurrence frequencies. Each patient contributes $(T_{i}^{(1)},T_{i}^{(2)})$-length multimodal data, so the variance of empirical co-occurrence frequencies is $O(1/(n(S_{1}^{(1)}\wedge S_{1}^{(2)})))$. Normalizing to obtain PMI entries scales this variance by $d_{1}d_{2}$, and converting to Frobenius norm adds an extra factor of $d_{1}^{1/2}d_{2}^{1/2}$ via the inequality $\|A\|_{\mathrm{F}}\leq d_{1}^{1/2}d_{2}^{1/2}\|A\|_{\max}$. Applying the Davis–Kahan theorem then gives a contribution of order $O(pd_{1}^{3/2}d_{2}^{3/2}/(n(S_{1}^{(1)}\wedge S_{1}^{(2)})))$. The second variance term reflects the averaging across $n$ patients, yielding a $1/\sqrt{n}$ rate, which again is scaled by $d_{1}^{1/2}d_{2}^{1/2}$ when converted to Frobenius norm. As a result, to make the overall bound vanish and ensure consistency, Assumption 3.2 is required.
3.2 Comparison with the baseline methods
We now compare CLAIMEwith the two baseline approaches introduced in Section 2.3. As discussed there, both baselines can be interpreted as applying SVD to different empirical PMI matrices: the first is SVD on the concatenated PMI matrix $\widehat{\mathbb{PMI}}$, while the second is based on the estimator $\widehat{\mathbb{PMI}}_{\textnormal{CL}}$ derived from the CL loss (2.5). We show in Theorem 3.4 (proof in Supplementary S4.4) that both $\widehat{\mathbb{PMI}}$ and $\widehat{\mathbb{PMI}}_{\textnormal{CL}}$ are biased estimators of the population matrix $\mathbb{PMI}$. Building on this result, Theorem 3.5 establishes the convergence rates of the corresponding estimators $\widehat{\mathbf{V}}_{\textnormal{CL}}$ and $\mathbf{\widehat{V}}_{\textnormal{Con}}$, with detailed proofs provided in Supplementary S4.4.
@@ 3.2 Comparison with the baseline methods
Under Assumptions 3.1-3.3 and 3.4, with probability $1-\exp(-\Omega(\log^{2}(n+d)))$,
$\norm{\widehat{\mathbb{PMI}}-(\mathbb{PMI}+\mathbf{B}_{\textnormal{Con}})}_{\mathrm{F}}$ $\lesssim\frac{(d_{1}\vee d_{2})d_{1}^{2}\log(n+d)}{\sqrt{n}S_{1}^{(1)1/2}}+\frac{(d_{1}\vee d_{2})d_{2}^{2}\log(n+d)}{\sqrt{n}S_{1}^{(2)1/2}}$
$\quad+\frac{(d_{1}\vee d_{2})d_{1}d_{2}\log(n+d)}{\sqrt{n}(S_{1}^{(1)1/2}\wedge S_{1}^{(2)1/2})}+\frac{(d_{1}\vee d_{2})\log(n+d)}{\sqrt{n}},\quad\text{ and}$
$\norm{\widehat{\mathbb{PMI}}_{\textnormal{CL}}-(\mathbb{PMI}+\mathbf{B}_{\textnormal{CL}})}_{\mathrm{F}}\lesssim\frac{(d_{1}\vee d_{2})d_{1}^{2}\log(n+d)}{\sqrt{n}S_{1}^{(1)1/2}}+\frac{(d_{1}\vee d_{2})d_{2}^{2}\log(n+d)}{\sqrt{n}S_{1}^{(2)1/2}}$
$\quad\quad+\frac{(d_{1}\vee d_{2})d_{1}d_{2}\log(n+d)}{\sqrt{n}(S_{1}^{(1)1/2}\wedge S_{1}^{(2)1/2})}+\frac{\log(n+d)(d_{1}\vee d_{2})}{\sqrt{n}}+\frac{p^{2}(d_{1}\vee d_{2})\log(n+d)}{(d_{1}\wedge d_{2})^{2}}.$
Here, $\mathbf{B}_{\textnormal{Con}}$ and $\mathbf{B}_{\textnormal{CL}}$ are bias matrices defined in Supplementary S4.2 and S4.4.
Remark 3.2.
As shown in Theorems S4.2 (Concate) and S4.5 (CL), the bias matrix $\mathbf{B}_{\textnormal{CL}}$ includes a correctable term based on observed $T_{i}^{(M)}$ and an uncorrectable term involving the population-level PMI. In contrast, $\mathbf{B}_{\textnormal{Con}}$ depends only on $T_{i}^{(M)}$ and is fully correctable. To evaluate the effect of this correction, we conducted additional experiments by removing the bias term associated with $T_{i}^{(M)}$. The results, presented in Supplementary S7.3.3, show that the debiased estimators offer little to no improvement over the original ones, suggesting that the impact of the removable bias component is limited in practice.
@@
Suppose that $\|\mathbf{B}_{\textnormal{Con}}\|\vee\|\mathbf{B}_{\textnormal{CL}}\|\ll 1/p$. Under Assumptions 3.1-3.4, with probability $1-\exp(-\Omega(\log^{2}(n+d)))$,
$\norm{\sin\Theta\quantity(\mathbf{P}_{p}(\mathbf{\widehat{V}}_{\textnormal{Con}}^{\scriptscriptstyle\sf T}),\mathbf{P}_{p}\quantity(\frac{1}{p}\mathbf{V}^{\star}\mathbf{V}^{\star{\scriptscriptstyle\sf T}}+{\bm{\Sigma}}+\mathbf{B}_{\textnormal{Con}}))}_{\mathrm{F}}$
$\quad\lesssim\frac{p(d_{1}\vee d_{2})d_{1}^{2}\log(n+d)}{\sqrt{n}S_{1}^{(1)1/2}}+\frac{p(d_{1}\vee d_{2})d_{2}^{2}\log(n+d)}{\sqrt{n}S_{1}^{(2)1/2}}$
$\quad\quad+\frac{p(d_{1}\vee d_{2})d_{1}d_{2}\log(n+d)}{\sqrt{n}(S_{1}^{(1)1/2}\wedge S_{1}^{(2)1/2})}+\frac{p\log(n+d)(d_{1}\vee d_{2})}{\sqrt{n}}+\frac{p^{3}(d_{1}\vee d_{2})\log(n+d)}{(d_{1}\wedge d_{2})^{2}},$
$\norm{\sin\Theta\quantity(\mathbf{P}_{p}(\widehat{\mathbf{V}}_{\textnormal{CL}}^{\scriptscriptstyle\sf T}),\mathbf{P}_{p}\quantity(\frac{1}{p}\mathbf{V}^{\star}\mathbf{V}^{\star{\scriptscriptstyle\sf T}}+{\bm{\Sigma}}+\mathbf{B}_{\textnormal{CL}}))}_{\mathrm{F}}$
$\quad\lesssim\frac{p(d_{1}\vee d_{2})d_{1}^{2}\log(n+d)}{\sqrt{n}S_{1}^{(1)1/2}}+\frac{p(d_{1}\vee d_{2})d_{2}^{2}\log(n+d)}{\sqrt{n}S_{1}^{(2)1/2}}$
$\quad\quad+\frac{p(d_{1}\vee d_{2})d_{1}d_{2}\log(n+d)}{\sqrt{n}(S_{1}^{(1)1/2}\wedge S_{1}^{(2)1/2})}+\frac{p\log(n+d)(d_{1}\vee d_{2})}{\sqrt{n}}+\frac{p^{3}(d_{1}\vee d_{2})\log(n+d)}{(d_{1}\wedge d_{2})^{2}}.$
Theorem 3.5 shows that both Concate and CL will yield biased estimators for the embeddings $\mathbf{V}^{\star}$ due to the inherent bias terms and covariance matrix of the noise. While the bias terms $\mathbf{B}_{\textnormal{Con}}$ and $\mathbf{B}_{\textnormal{CL}}$ can be partially removed, the influence of $\bm{\Sigma}$} remains irreducible due to its unobserved nature. Comparing Theorems 3.3 and 3.5, we find that CLAIME achieves a sharper rate than Concate and CL even when $\mathbf{B}_{\textnormal{CL}}$, $\mathbf{B}_{\textnormal{Con}}$ and $\bm{\Sigma}$} go to zero. As a result, CLAIME is a better estimator than Concate and CL. For the bias terms $\mathbf{B}_{\textnormal{Con}}$ and $\mathbf{B}_{\textnormal{CL}}$, we have the following lemma. The proof of Lemma 3.1 is given in Supplementary S4.7.
Lemma 3.1.
Suppose that Assumption 3.4 holds. If $T_{i}^{(1)}=T_{i}^{(2)}$, then, $\|\mathbf{B}_{\textnormal{Con}}\|\vee\|\mathbf{B}_{\textnormal{CL}}\|\lesssim(d_{1}\vee d_{2})S_{1}^{(1)}/S_{2}^{(1)2}$. Furthermore, if $T_{i}^{(1)}=T_{i}^{(2)}\equiv T$ for all $i\in[n]$ for some $T>0$, then $\|\mathbf{B}_{\textnormal{Con}}\|\vee\|\mathbf{B}_{\textnormal{CL}}\|\lesssim(d_{1}\vee d_{2})/T$.
We then characterize the effect of $\bm{\Sigma}$} and show that when $\bm{\Sigma}$} satisfies suitable conditions, the two existing methods Concate and CL will not provide ideal estimators of $\mathbf{V}^{\star}$.
@@
Assume that $(s_{p}({\bm{\Sigma}})-s_{p+1}({\bm{\Sigma}}))/s_{p}({\bm{\Sigma}})\gtrsim 1$.
Assumption 3.5 is known as the eigengap condition for the top-$p$ singular vectors of $\bm{\Sigma}$}. A similar assumption has been used in the analysis of contrastive learning (Ji et al., , 2023). This enables us to identify the top-$p$ singular vectors of $(1/p)\mathbf{V}^{\star}\mathbf{V}^{\star{\scriptscriptstyle\sf T}}+{\bm{\Sigma}}$.
@@
Suppose $\|\mathbf{B}_{\textnormal{Con}}\|\vee\|\mathbf{B}_{\textnormal{CL}}\|\ll 1/p$. Under Assumptions 3.1-3.6, with probability $1-\exp(-\Omega(\log^{2}(n+d)))$, $\|\sin\Theta(\mathbf{\widehat{U}}_{\textnormal{Con}},\mathbf{U}^{\star})\|_{\mathrm{F}}\gtrsim\sqrt{p}$ and $\|\sin\Theta(\mathbf{\widehat{U}}_{\textnormal{CL}},\mathbf{U}^{\star})\|_{\mathrm{F}}\gtrsim\sqrt{p}$.
@@
Choose $\mathbf{P}^{(1)}\in\mathcal{O}_{d_{1},p}$ satisfying $\mathbf{U}_{1}^{\star};\mathbf{P}^{(1)}]\in\mathcal{O}_{d_{1},2p}$ and $\|[\mathbf{U}_{1}^{\star};\mathbf{P}^{(1)}]\|_{2,\infty}^{2}\lesssim p/d_{1}$. Similarly choose $\mathbf{P}^{(2)}\in\mathcal{O}_{d_{2},p}$ such that $\mathbf{U}_{2}^{\star};\mathbf{P}^{(2)}]\in\mathcal{O}_{d_{2},2p}$ and $\|[\mathbf{U}_{2}^{\star};\mathbf{P}^{(2)}]\|_{2,\infty}^{2}\lesssim p/d_{2}$. Let ${\bm{\Sigma}}_{1}:=\mathbf{P}^{(1)}\mathbf{P}^{(1){\scriptscriptstyle\sf T}}$, ${\bm{\Sigma}}_{2}:=(1/2)\mathbf{P}^{(2)}\mathbf{P}^{(2){\scriptscriptstyle\sf T}}$, $\mathbf{V}^{\star}_{1}=\mathbf{U}^{\star}_{1}$ and $\mathbf{V}^{\star}_{2}=\mathbf{U}^{\star}_{2}$. For these noise covariance matrices, Assumptions 3.1, 3.3 and 3.5 hold. In addition, Assumption 3.6 holds since $\mathbf{P}^{(1){\scriptscriptstyle\sf T}}\mathbf{V}_{1}^{\star}=\mathbf{P}^{(2){\scriptscriptstyle\sf T}}\mathbf{V}_{2}^{\star}=\mathbf{O}_{p\times p}$. Under additional assumptions on $(T_{i}^{(M)})_{i\in[n]}$, $d_{1}$ and $d_{2}$ (Assumptions 3.2 and 3.4), with probability $1-\exp(-\Omega(\log^{2}(n+d)))$,
@@
$\min_{\mathbf{H}\in\mathcal{O}_{p,p}}\|\sin\Theta(\mathbf{\widehat{U}}_{\mathbf{H}},\mathbf{U}^{\star})\|_{\mathrm{F}}$ $\lesssim\left\{\frac{p^{3}(d_{1}\vee d_{2})}{(d_{1}\wedge d_{2})^{2}}+\frac{pd_{1}^{3/2}d_{2}^{3/2}}{\sqrt{n}(S_{1}^{(1)1/2}\wedge S_{1}^{(2)1/2})}+\frac{pd_{1}^{1/2}d_{2}^{1/2}}{\sqrt{n}}\right\}\log(n+d).$
Corollary 3.1 directly follows from Theorems 3.3 and 3.6. Note that the factor $1/2$ in the definition of ${\bm{\Sigma}}_{2}$ guarantees the eigengap condition at the $p$-th largest singular value of the joint noise covariance matrix $\bm{\Sigma}$}. The noise covariance in Corollary 3.1 implies that the subspace spanned by the noise is orthogonal to that of the signal, so that noisy fluctuations in the feature frequency distribution occur simultaneously across many features, in a way separable from the signal. Under this condition, Concate and CL tend to learn the eigenspace corresponding to noise covariance matrix, as they are vulnerable to strong noise. In contrast, CLAIME effectively escapes from the noise and recovers the core representations when $n$ and $(d_{1}\vee d_{2})/(d_{1}\wedge d_{2})^{2}$ are sufficiently large.